OpenStudy (bethgrace96):

Please answer! MEDAL! three consecutive integers are such that the sum of the first integer and the product of the other two is 62. which equation could be used to find the integers? x^2-x-1=62 2x^2-1=62 x^2+4x+2=62 2x^2+X+1=62

OpenStudy (anonymous):

You need to set this up as a formula with (x) being the smallest integer.

OpenStudy (anonymous):

x^2+4x+2=62 If the numbers are expressed as follows, {x - 1, x, x + 1} then x^2+2 x-1=62

OpenStudy (bethgrace96):

so whats the answer?

OpenStudy (anonymous):

I think it needs to be expressed as (x,x+1,x+2) In that the problem states x is the first integer.

OpenStudy (bethgrace96):

so whats the answer.........

OpenStudy (anonymous):

x^2+4x+2=62

OpenStudy (anonymous):

You have to write out the equation (x) + (x+1(x+2)